[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Apr 03, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Apr 3 19:58:57 UTC 2017


HSU Shift Log for Mon Apr 03 2017 15:58:50 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: 12Z GFS data still not totally in at 4pm, with a big gap  
from Day 7-10. Some serious data transfer issues for this cycle.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170403/62f68710/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list